OpenStudy (anonymous):

Create your own unique quadratic equation in the form y = ax2 + bx + c. Use complete sentences and show all work to determine the following: Does the graph open up or down? How do you know? Explain whether the graph has a maximum or minimum point. Find the vertex and x-intercepts of the graph.

OpenStudy (nincompoop):

review this example: http://openstudy.com/study#/updates/516b8bb2e4b02ec89c5ab7c2

OpenStudy (anonymous):

Im really confused

OpenStudy (nincompoop):

CLICK ON THE LINK :)

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

so if i made the equation 20x^2+15x+10 how do I get the vertex and x intercepts?

OpenStudy (nincompoop):

so that example is a quadratic function in that tutorial, I went in great detail in determining whether it opened upward or downward. as for the form of a quadratic function: y=ax^2+bx+c a quadratic equation has its first term with an exponent of 2

OpenStudy (anonymous):

ya?

OpenStudy (nincompoop):

read everything in the link I provided you will learn how to get the vertex of a quadratic function

OpenStudy (anonymous):

i have the notes written down I just forgot how to do it

OpenStudy (nincompoop):

I indicated how to look for vertex in the link I provided. step-by-step

OpenStudy (anonymous):

okay so if I use 25x^2+15x+10 the graph is going up because its positive idk idk

OpenStudy (anonymous):

okay sooo i use x=-b/2a to solvethe vertex

OpenStudy (anonymous):

how do I solve the x-intecepts?

OpenStudy (anonymous):

this is hard!!!!!!!!!!!!!!

OpenStudy (anonymous):

vertex= 15/10(-2) = 15/-20 = 0.17 -_-

OpenStudy (anonymous):

0.75 i meant

OpenStudy (nincompoop):

well you have to understand the whole concept and this takes an explanation. and it is just about using -b/2a

OpenStudy (anonymous):

the x-intecept is to?

OpenStudy (nincompoop):

you will find out what vertex is and that you need not worry about separately solving for x-intercept if you read carefully...

OpenStudy (anonymous):

the vertex is 1/2?

OpenStudy (nincompoop):

no... please read

OpenStudy (nincompoop):

@mathslover

mathslover (mathslover):

Yes?

OpenStudy (nincompoop):

help the girl out